In a PMDC motor, when the armature is rotated at 100rpm, an emf of 2V is developed across its terminals. For this motor, calculate the armature current flow required to generate a torque of 0.4Nm


1
Expert's answer
2020-10-25T19:47:34-0400

Required torque("T" ) = 0.4 Nm

"Eb" =2V

Mechanical power delivered ="P = Eb*Ia =T * \omega"

Where "Eb" = Back emf and , "Ia" = Armature current, "\omega" = anguar frequency.


"\omega = 2\pi*n/60" "= 2\pi*100/60" = 10.472 rad/sec

Therefore Ia = "T*\omega/Eb" = 0.4*10.472/2 = 2.0944 A